# Hodge diamonds of the Landau–Ginzburg orbifolds

### 26 Feb 2024 v2

@@ 2.2. Graph of a quasihomogeneous singularity
Figure 1. Graphs of $N=3$ quasihomogenous singularities (see Example 4)
@@
Figure 2. Connected component of $\Gamma_{f}$ (see Proposition 6 )
It’s easy to see that $\Gamma_{f\oplus g}=\Gamma_{f}\sqcup\Gamma_{g}$, but it’s not true that $f$ decomposes into the $\oplus$–sum if $\Gamma_{f}$ has more than one component. For example the graph of $f=x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{1}x_{2}x_{3}$ is just the disjoint union of three vertices without any edges.
2.3. Graph decomposition of a polynomial
Assume we only know the graph $\Gamma_{f}$ and not the polynomial $f$ itself. The graph structure indicates some monomials that are the summands of $f$. Call these monomials graph monomials. In particular, $f$ has only graph monomials if it is of Fermat, chain or loop type or a $\oplus$–sum of them.
@@ 2.3. Graph decomposition of a polynomial
    $f_{0},f_{1},\dots,f_{p},f_{add}\in{\mathbb{C}}[x_{1},\dots,x_{N}]$,
@@
    $f_{add}:=f-f_{0}-f_{1}\dots-f_{p}$ having as the summands all the non–graph monomials of $f$.
@@
Example 7.
The polynomial $f=x_{1}^{3}+x_{1}\left(x_{2}^{2}+x_{3}^{2}+x_{4}^{2}\right)+\epsilon x_{2}x_{3}x_{4}$ with $\varepsilon\in{\mathbb{C}}\backslash\{0,\pm 2\sqrt{-1}\}$ defines an isolated singularity. It’s also quasihomogeneous with $q_{1}=\dots=q_{4}=1/3$. We have $p=3$,
$f_{0}=x_{1}^{3},\ f_{1}=x_{1}x_{2}^{2},\ f_{2}=x_{1}x_{3}^{2},\ f_{3}=x_{1}x_{4}^{2},\ f_{add}=\varepsilon x_{2}x_{3}x_{4}.$
2.4. Graph exponents matrix
Let $f$ define a quasihomogeneous singularity. We introduce the matrix $E_{f}$ with the entries in ${\mathbb{Z}}_{\geq 0}$. It follows from Proposition 6 that $f$ has exactly $N$ graph monomials. Let every row of $E_{f}$ correspond to a graph monomial. The components of this row will be $(\alpha_{1},\dots,\alpha_{N})$ if and only if the corresponding graph monomial is $\varepsilon\cdot x_{1}^{\alpha_{1}}\cdots x_{N}^{\alpha_{N}}$ for some $\varepsilon\in{\mathbb{C}}^{\ast}$.
@@ 2.4. Graph exponents matrix
Let $E_{ij}$ denote the components of $E_{f}$. Then for some non–zero constants $c_{k}$ we have
$f-f_{add}=\sum_{k=1}^{N}c_{k}x_{1}^{E_{k1}}\cdots x_{N}^{E_{kN}}.$ (5)
Remark 8.
@@
Example 9.
The matrices $E_{f}$ of the Example 4 are
$E_{f_{I}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&0\\ 0&0&a_{3}\end{pmatrix},\ E_{f_{II}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&1\\ 0&0&a_{3}\end{pmatrix},\ E_{f_{III}}=\begin{pmatrix}a_{1}&0&0\\ 1&a_{2}&0\\ 1&0&a_{3}\end{pmatrix},\ E_{f_{IV}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&1\\ 1&0&a_{3}\end{pmatrix},$
$E_{f_{V}}=\begin{pmatrix}a_{1}&0&0\\ 1&a_{2}&0\\ 0&1&a_{3}\end{pmatrix},\ E_{f_{VI}}=\begin{pmatrix}a_{1}&1&0\\ 1&a_{2}&0\\ 1&0&a_{3}\end{pmatrix},\ E_{f_{VII}}=\begin{pmatrix}a_{1}&1&0\\ 0&a_{2}&1\\ 1&0&a_{3}\end{pmatrix}.$
@@
$E_{loop}=\left(\begin{array}[]{ccccc}a_{1}&1&\dots&0&0\\ 0&a_{2}&1&\dots&0\\ \vdots&&\ddots&\ddots&0\\ 0&0&\dots&a_{N-1}&1\\ 1&0&0&\dots&a_{N}\end{array}\right),\quad E_{chain}=\left(\begin{array}[]{ccccc}a_{1}&0&\dots&0&0\\ 1&a_{2}&0&\dots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\dots&1&a_{N-1}&0\\ 0&0&\dots&1&a_{N}\end{array}\right)$ (6)
In general, by Proposition 6 if $\Gamma_{f}$ has only one connected component, the matrix $E_{f}$ after some renumbering of the variables and the rows has the block form. The diagonal blocks are several chain type exponent matrices and exactly one loop type exponents matrix as in Eq. (6), such that for every chain type block there is exactly one additional matrix entry $1$ in the first row of this block and the column of a loop type block. All the other matrix entries except listed vanish.
$E_{f}=\left(\begin{array}[]{cccc}\hfil A_{0}\hfil\vrule&0&0&0\\ \hline\cr\hfil U_{i_{1}j_{1}}\hfil\vrule&\hfil A_{1}\hfil\vrule&0&0\\ \hline\cr\vdots&&\ddots&0\\ \hline\cr\hfil U_{i_{p}j_{p}}\hfil\vrule&0&0&\vrule\hfil A_{p}\hfil\vrule\end{array}\right)$ (7)
@@
Assuming the decomposition of Eq. (4), the matrix $A_{0}$ is exactly the exponent matrix of $f_{0}$ and the matrices $A_{1},\dots,A_{p}$ are defined by $f_{1},\dots,f_{p}$.
Proposition 10.
@@
Let $f$ be decomposed as in Eq. (4). Then $f_{0}$ is of Fermat or loop type and the weights of its variables are defined uniquely. Similarly for any $f_{k}$ with $k=1,\dots,p$ corresponding to the tree with the root on the oriented circle, the weight of the root’s variable is defined by the quasihomogeneity of $f_{0}$, going up the tree of deduces uniquely the weight of every variable of $f_{k}$ corresponding to the consequent vertex.
Introduce two ${\mathbb{Z}}^{N}$ vectors: $\overline{q}:=(q_{1},\dots,q_{N})^{T}$ and $\boldsymbol{1}:=(1,\dots,1)^{T}$. Then the quasihomogeneity condition on $f$ is equivalent to the ${\mathbb{Q}}^{N}$ vector equality $E_{f}\cdot\overline{q}=\boldsymbol{1}$. It follows now from Cramer rule that $\det(E_{f})\neq 0$ because this equation has a unique solution. This completes (i).
The canonical weight set is obtained by taking $d_{f}:=\det(E_{f})$ and solving $E_{f}\cdot\overline{d}=d_{f}\boldsymbol{1}$ for $\overline{d}:=(d_{1},\dots,d_{N})^{T}$. ∎
2.5. Invertible polynomials
@@
The following statement can be assumed as a well-known (cf. [KS]), we add up the proof for completeness.
Proposition 11.
Let $f$ be an invertible polynomial. Then after some rescaling and renumbering of the variables we have $f=f^{(1)}\oplus\dots\oplus f^{(n)}$ for $f^{(k)}$ being either of Fermat, chain or loop type.
@@
$\frac{\partial f}{\partial x_{i}}$ $=a\alpha_{1}x_{i}^{a-1}x_{l}^{K}+\alpha_{2}x_{j}^{b}+\alpha_{3}x_{k}^{c},$ (8)
$\frac{\partial f}{\partial x_{j}}$ $=b\alpha_{2}x_{j}^{b-1}x_{i},\ \frac{\partial f}{\partial x_{k}}=a\alpha_{3}x_{k}^{c-1}x_{i}.$ (9)
Setting $x_{i}=0$ we see that vanishing $\frac{\partial f}{\partial x_{i}}=\frac{\partial f}{\partial x_{j}}=\frac{\partial f}{\partial x_{k}}=0$ is equivalent to $\alpha_{2}x_{j}^{b}+\alpha_{3}x_{k}^{c}=0$ what what shows that $x_{i}=x_{j}=x_{k}=0$ is not an isolated critical point of $f$. ∎
@@
In the notation of Section 2.3 the graph decomposition Eq. (4) of Fermat, loop and chain type polynomials is the following. We have always $f_{add}=0$, $p=0$ and $f_{0}=f$ for Fermat and loop types, but $p=1$ and $f_{0}+f_{1}=f$ for chain type with $f_{0}=x_{m}^{a_{m}}$.
Example 12.
@@ 3. Symmetries
${\mathrm{GL}}_{f}:=\left\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\ |\ f(g\cdot{\bf x})=f({\bf x})\right\}.$
Lemma 13.
Under our assumptions on $f$ any $g\in{\mathrm{GL}}_{f}$ necessarily preserves the weights of the variables, i.e. maps each $x_{i}$ to a linear combination of $x_{j}$ with the same weight.
@@
We may now assume that $f$ has no quadratic terms. In this case, the argument of [OS78, Theorem 2.1] applies verbatim to the quasihomogeneous situation to prove that ${\mathrm{GL}}_{f}$ is finite. Let
$\widetilde{{\mathrm{GL}}}_{f}=\left\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\ |\ f(g\cdot{\bf x})=\chi(g)f({\bf x}),\chi(g)\in{\mathbb{C}}^{*}\right\}.$
The map $\chi\colon\widetilde{{\mathrm{GL}}}_{f}\to{\mathbb{C}}^{*}$ is a character and ${\mathrm{GL}}_{f}$ is precisely the kernel of $\chi$, in particular, it is a normal subgroup. Moreover, the condition (QH1) provides an inclusion $t\colon{\mathbb{C}}^{*}\hookrightarrow\widetilde{{\mathrm{GL}}}_{f}$, such that $\chi\circ t$ is a degree $d_{0}>0$ map. The action by conjugation of connected subgroup $t({\mathbb{C}}^{*})$ on the finite subgroup ${\mathrm{GL}}_{f}$ is necessary trivial, which means that ${\mathrm{GL}}_{f}$ commute with $t({\mathbb{C}}^{*})$. This means that ${\mathrm{GL}}_{f}$ preserves the eigenspaces of $t({\mathbb{C}}^{*})$ as desired. ∎
Remark 14.
Notably, the observation of this Lemma seems to be unknown to previous authors (for example, it was imposed as condition in [M] and some subsequent works).
@@
$G_{f}^{d}\cong\left\{(\lambda_{1},\dots,\lambda_{N})\in({\mathbb{C}}^{*})^{N}\ |\ f(\lambda_{1}x_{1},\dots,\lambda_{N}x_{N})=f(x_{1},\dots,x_{N})\right\}.$
@@
$G_{f^{\prime}\oplus f^{\prime\prime}}^{d}\cong G_{f^{\prime}}^{d}\times G_{f^{\prime\prime}}^{d}.$ (10)
@@
${\bf e}\left[\alpha\right]:=\exp(2\pi\sqrt{-1}\alpha),\quad\alpha\in{\mathbb{R}}.$
@@
$g={\rm diag}\left({\bf e}\left[\frac{\alpha_{1}}{r}\right],\dots,{\bf e}\left[\frac{\alpha_{N}}{r}\right]\right)\quad\mbox{with }0\leq\alpha_{i}<r,\alpha_{i}\in{\mathbb{N}}$ (11)
@@
${\overline{g}=(\alpha_{1}/r,\dots,\alpha_{N}/r)}\quad\text{ or }\quad{\overline{g}=\frac{1}{r}(\alpha_{1},\dots,\alpha_{N})}$
@@
Example 15.
For $f=x_{1}^{a_{1}}$ we have ${\mathrm{GL}}_{f}=G_{f}^{d}=\langle g\rangle$ with $g\in{\mathbb{C}}^{\ast}$ acting by $g(x_{1})={\bf e}\left[\frac{1}{a_{1}}\right]\cdot x_{1}$. Its order is $a_{1}$ and in the additive notation we have $\overline{g}=(1/a_{1})$, giving us ${\mathrm{GL}}_{f}\cong{\mathbb{Z}}/a_{1}{\mathbb{Z}}$.
Example 16.
For $f=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}$ we have $G_{f}^{d}=\langle g_{1},g_{2}\rangle$ with $g_{1}\cdot(x_{1},x_{2})=({\bf e}[\frac{1}{a_{1}}]x_{1},x_{2})$ and ${g_{2}\cdot(x_{1},x_{2})=({\bf e}[\frac{1}{a_{2}}(1-\frac{1}{a_{1}})]x_{1},{\bf e}[\frac{1}{a_{2}}]x_{2})}$. In the additive notation $\overline{g}_{1}=(1/a_{1},0)$ and $\overline{g}_{2}=((a_{1}-1)/(a_{1}a_{2}),1/a_{2})$. In this example ${\mathrm{GL}}_{f}=G_{f}^{d}$ because $q_{1}\neq q_{2}$.
@@
In particular it follows that $G_{f}^{d}$ and ${\mathrm{GL}}_{f}$ are not empty whenever $f$ is quasihomogeneous. Denote by $J$ the group generated by $j_{f}$:
@@
3.1. Fixed loci of the ${\mathrm{GL}}_{f}$ elements

### 25 Mar 2024 v3

@@ 2.2 Graph of a quasihomogeneous singularity
Figure 1: Graphs of $N=3$ quasihomogenous singularities (see Example 2.3).
@@ 2. $(2)$
Figure 2: Connected component of $\Gamma_{f}$ (see Proposition 2.5).
It’s easy to see that $\Gamma_{f\oplus g}=\Gamma_{f}\sqcup\Gamma_{g}$, but it is not true that $f$ decomposes into the $\oplus$-sum if $\Gamma_{f}$ has more than one component. For example the graph of $f=x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{1}x_{2}x_{3}$ is just the disjoint union of three vertices without any edges.
2.3 Graph decomposition of a polynomial
Assume we only know the graph $\Gamma_{f}$ and not the polynomial $f$ itself. The graph structure indicates some monomials that are the summands of $f$. Call these monomials graph monomials. In particular, $f$ has only graph monomials if it is of Fermat, chain or loop type or a $\oplus$-sum of them.
@@ 2.3 Graph decomposition of a polynomial
    $f_{0},f_{1},\dots,f_{p},f_{\rm add}\in{\mathbb{C}}[x_{1},\dots,x_{N}]$,
@@
    $f_{\rm add}:=f-f_{0}-f_{1}-\dots-f_{p}$ having as the summands all the non-graph monomials of $f$.
@@
Example 2.6.
The polynomial $f=x_{1}^{3}+x_{1}\left(x_{2}^{2}+x_{3}^{2}+x_{4}^{2}\right)+\epsilon x_{2}x_{3}x_{4}$ with $\varepsilon\in{\mathbb{C}}\backslash\left\{0,\pm 2\sqrt{-1}\right\}$ defines an isolated singularity. It’s also quasihomogeneous with $q_{1}=\dots=q_{4}=1/3$. We have $p=3$,
$f_{0}=x_{1}^{3},\qquad f_{1}=x_{1}x_{2}^{2},\qquad f_{2}=x_{1}x_{3}^{2},\qquad f_{3}=x_{1}x_{4}^{2},\qquad f_{\rm add}=\varepsilon x_{2}x_{3}x_{4}.$
2.4 Graph exponents matrix
Let $f$ define a quasihomogeneous singularity. We introduce the matrix $E_{f}$ with the entries in ${\mathbb{Z}}_{\geq 0}$. It follows from Proposition 2.5 that $f$ has exactly $N$ graph monomials. Let every row of $E_{f}$ correspond to a graph monomial. The components of this row will be $(\alpha_{1},\dots,\alpha_{N})$ if and only if the corresponding graph monomial is $\varepsilon\cdot x_{1}^{\alpha_{1}}\cdots x_{N}^{\alpha_{N}}$ for some $\varepsilon\in{\mathbb{C}}^{\ast}$.
@@ 2.4 Graph exponents matrix
Let $E_{ij}$ denote the components of $E_{f}$. Then for some non-zero constants $c_{k}$ we have
$f-f_{\rm add}=\sum_{k=1}^{N}c_{k}x_{1}^{E_{k1}}\cdots x_{N}^{E_{kN}}.$ (2.3)
Remark 2.7.
@@
Example 2.8.
The matrices $E_{f}$ of the Example 2.3 are
$E_{f_{\rm I}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&0\\ 0&0&a_{3}\end{pmatrix},\qquad$ $E_{f_{\rm II}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&1\\ 0&0&a_{3}\end{pmatrix},\qquad$ $E_{f_{\rm III}}=\begin{pmatrix}a_{1}&0&0\\ 1&a_{2}&0\\ 1&0&a_{3}\end{pmatrix},$
$E_{f_{\rm IV}}=\begin{pmatrix}a_{1}&0&0\\ 0&a_{2}&1\\ 1&0&a_{3}\end{pmatrix},\qquad$ $E_{f_{\rm V}}=\begin{pmatrix}a_{1}&0&0\\ 1&a_{2}&0\\ 0&1&a_{3}\end{pmatrix},\qquad$ $E_{f_{\rm VI}}=\begin{pmatrix}a_{1}&1&0\\ 1&a_{2}&0\\ 1&0&a_{3}\end{pmatrix},$
$E_{f_{\rm VII}}=\begin{pmatrix}a_{1}&1&0\\ 0&a_{2}&1\\ 1&0&a_{3}\end{pmatrix}.$
@@
$E_{\rm loop}=\left(\begin{matrix}a_{1}&1&\dots&0&0\\ 0&a_{2}&1&\dots&0\\ \vdots&&\ddots&\ddots&0\\ 0&0&\dots&a_{N-1}&1\\ 1&0&0&\dots&a_{N}\end{matrix}\right),\qquad E_{\rm chain}=\left(\begin{matrix}a_{1}&0&\dots&0&0\\ 1&a_{2}&0&\dots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\dots&1&a_{N-1}&0\\ 0&0&\dots&1&a_{N}\end{matrix}\right).$ (2.4)
In general, by Proposition 2.5 if $\Gamma_{f}$ has only one connected component, the matrix $E_{f}$ after some renumbering of the variables and the rows has the block form. The diagonal blocks are several chain type exponent matrices and exactly one loop type exponents matrix as in equation (2.4), such that for every chain type block there is exactly one additional matrix entry $1$ in the first row of this block and the column of a loop type block. All the other matrix entries except listed vanish,
$E_{f}=\left(\begin{array}[]{cccc}\hfil A_{0}\hfil\vrule&0&0&0\\ \hline\cr\hfil U_{i_{1}j_{1}}\hfil\vrule&\hfil A_{1}\hfil\vrule&0&0\\ \hline\cr\vdots&&\ddots&0\\ \hline\cr\hfil U_{i_{p}j_{p}}\hfil\vrule&0&0&\vrule\hfil A_{p}\hfil\vrule\end{array}\,\right),$ (2.5)
@@
Assuming the decomposition of equation (2.2), the matrix $A_{0}$ is exactly the exponent matrix of $f_{0}$ and the matrices $A_{1},\dots,A_{p}$ are defined by $f_{1},\dots,f_{p}$.
Proposition 2.9.
@@ 2. $(ii)$
Let $f$ be decomposed as in equation (2.2). Then $f_{0}$ is of Fermat or loop type and the weights of its variables are defined uniquely. Similarly for any $f_{k}$ with $k=1,\dots,p$ corresponding to the tree with the root on the oriented circle, the weight of the root’s variable is defined by the quasihomogeneity of $f_{0}$, going up the tree of deduces uniquely the weight of every variable of $f_{k}$ corresponding to the consequent vertex.
Introduce two ${\mathbb{Z}}^{N}$ vectors: $\overline{q}:=(q_{1},\dots,q_{N})^{\mathsf{T}}$ and $\boldsymbol{1}:=(1,\dots,1)^{\mathsf{T}}$. Then the quasihomogeneity condition on $f$ is equivalent to the ${\mathbb{Q}}^{N}$ vector equality $E_{f}\cdot\overline{q}=\boldsymbol{1}$. It follows now from Cramer rule that $\det(E_{f})\neq 0$ because this equation has a unique solution. This completes (i).
The canonical weight set is obtained by taking $d_{f}:=\det(E_{f})$ and solving $E_{f}\cdot\overline{d}=d_{f}\boldsymbol{1}$ for $\overline{d}:=(d_{1},\dots,d_{N})^{\mathsf{T}}$. ∎
2.5 Invertible polynomials
@@
The following statement can be assumed as a well-known (cf. [23]), we add up the proof for completeness.
Proposition 2.10.
Let $f$ be an invertible polynomial. Then after some rescaling and renumbering of the variables we have $f=f^{(1)}\oplus\cdots\oplus f^{(n)}$ for $f^{(k)}$ being either of Fermat, chain or loop type.
@@
$\frac{\partial f}{\partial x_{i}}=a\alpha_{1}x_{i}^{a-1}x_{l}^{K}+\alpha_{2}x_{j}^{b}+\alpha_{3}x_{k}^{c},\qquad\frac{\partial f}{\partial x_{j}}=b\alpha_{2}x_{j}^{b-1}x_{i},\qquad\frac{\partial f}{\partial x_{k}}=a\alpha_{3}x_{k}^{c-1}x_{i}.$
Setting $x_{i}=0$, we see that vanishing $\frac{\partial f}{\partial x_{i}}=\frac{\partial f}{\partial x_{j}}=\frac{\partial f}{\partial x_{k}}=0$ is equivalent to $\alpha_{2}x_{j}^{b}+\alpha_{3}x_{k}^{c}=0$ what shows that $x_{i}=x_{j}=x_{k}=0$ is not an isolated critical point of $f$. ∎
@@
In the notation of Section 2.3, the graph decomposition equation (2.2) of Fermat, loop and chain type polynomials is the following. We have always $f_{\rm add}=0$, $p=0$ and $f_{0}=f$ for Fermat and loop types, but $p=1$ and $f_{0}+f_{1}=f$ for chain type with $f_{0}=x_{m}^{a_{m}}$.
Example 2.11.
@@ 3 Symmetries
${\mathrm{GL}}_{f}:=\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\mid f(g\cdot{\bf x})=f({\bf x})\}.$
Lemma 3.1.
Under our assumptions on $f$ any $g\in{\mathrm{GL}}_{f}$ necessarily preserves the weights of the variables, i.e., maps each $x_{i}$ to a linear combination of $x_{j}$ with the same weight.
@@
We may now assume that $f$ has no quadratic terms. In this case, the argument of [26, Theorem 2.1] applies verbatim to the quasihomogeneous situation to prove that ${\mathrm{GL}}_{f}$ is finite. Let
$\widetilde{{\mathrm{GL}}}_{f}=\{g\in{\mathrm{GL}}(N,{\mathbb{C}})\mid f(g\cdot{\bf x})=\chi(g)f({\bf x}),\chi(g)\in{\mathbb{C}}^{*}\}.$
The map $\chi\colon\widetilde{{\mathrm{GL}}}_{f}\to{\mathbb{C}}^{*}$ is a character and ${\mathrm{GL}}_{f}$ is precisely the kernel of $\chi$, in particular, it is a normal subgroup. Moreover, the condition (2.1) provides an inclusion $t\colon{\mathbb{C}}^{*}\hookrightarrow\widetilde{{\mathrm{GL}}}_{f}$, such that $\chi\circ t$ is a degree $d_{0}>0$ map. The action by conjugation of connected subgroup $t({\mathbb{C}}^{*})$ on the finite subgroup ${\mathrm{GL}}_{f}$ is necessary trivial, which means that ${\mathrm{GL}}_{f}$ commute with $t({\mathbb{C}}^{*})$. This means that ${\mathrm{GL}}_{f}$ preserves the eigenspaces of $t({\mathbb{C}}^{*})$ as desired. ∎
Remark 3.2.
Notably, the observation of this Lemma seems to be unknown to previous authors (for example, it was imposed as condition in [25] and some subsequent works).
@@
$G_{f}^{d}\cong\bigl{\{}(\lambda_{1},\dots,\lambda_{N})\in({\mathbb{C}}^{*})^{N}\mid f(\lambda_{1}x_{1},\dots,\lambda_{N}x_{N})=f(x_{1},\dots,x_{N})\bigr{\}}.$
@@
$G_{f^{\prime}\oplus f^{\prime\prime}}^{d}\cong G_{f^{\prime}}^{d}\times G_{f^{\prime\prime}}^{d}.$ (3.1)
@@
${\bf e}\left[\alpha\right]:=\exp\bigl{(}2\pi\sqrt{-1}\alpha\bigr{)},\qquad\alpha\in{\mathbb{R}}.$
@@
$g={\rm diag}\left({\bf e}\left[\frac{\alpha_{1}}{r}\right],\dots,{\bf e}\left[\frac{\alpha_{N}}{r}\right]\right)\qquad\mbox{with}\quad 0\leq\alpha_{i}<r,\quad\alpha_{i}\in{\mathbb{N}},$
@@
${\overline{g}=(\alpha_{1}/r,\dots,\alpha_{N}/r)}\qquad\text{ or }\qquad{\overline{g}=\frac{1}{r}(\alpha_{1},\dots,\alpha_{N})}$
@@
Example 3.3.
For $f=x_{1}^{a_{1}}$ we have ${\mathrm{GL}}_{f}=G_{f}^{d}=\langle g\rangle$ with $g\in{\mathbb{C}}^{\ast}$ acting by $g(x_{1})={\bf e}\bigl{[}\frac{1}{a_{1}}\bigr{]}\cdot x_{1}$. Its order is $a_{1}$ and in the additive notation we have $\overline{g}=(1/a_{1})$, giving us ${\mathrm{GL}}_{f}\cong{\mathbb{Z}}/a_{1}{\mathbb{Z}}$.
Example 3.4.
For $f=x_{1}^{a_{1}}x_{2}+x_{2}^{a_{2}}$ we have $G_{f}^{d}=\langle g_{1},g_{2}\rangle$ with $g_{1}\cdot(x_{1},x_{2})=({\bf e}[\frac{1}{a_{1}}]x_{1},x_{2})$ and ${g_{2}\cdot(x_{1},x_{2})=\bigl{(}{\bf e}[\frac{1}{a_{2}}(1-\frac{1}{a_{1}})]x_{1},{\bf e}\bigl{[}\frac{1}{a_{2}}\bigr{]}x_{2}\bigr{)}}$. In the additive notation $\overline{g}_{1}=(1/a_{1},0)$ and $\overline{g}_{2}=((a_{1}-1)/(a_{1}a_{2}),1/a_{2})$. In this example ${\mathrm{GL}}_{f}=G_{f}^{d}$ because $q_{1}\neq q_{2}$.
@@
In particular, it follows that $G_{f}^{d}$ and ${\mathrm{GL}}_{f}$ are not empty whenever $f$ is quasihomogeneous. Denote by $J$ the group generated by $j_{f}$:
@@
3.1 Fixed loci of the $\boldsymbol{{\mathrm{GL}}_{f}}$ elements